# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 7 Exploiting integrability
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $\mathbb{B}$ be its right-angled building, and let $H$ be a countable group with bounded torsion. Assume that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$. Then $H$ acts on $\mathbb{B}$ with virtually infinite cyclic stabilizers of rank $1$ vertices.
@@ 7.1 Integrable coupling between vertex stabilizers
Let $G,H$ be countable groups, with $G$ finitely generated, and let $\hat{G}$ be a finite-index extension of $G$. Let $\hat{\Omega}$ be an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $\hat{G}$. Then $\hat{\Omega}$ is an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$.
@@
By definition, there exists a fundamental domain $X_{\hat{G}}$ for the action of $\hat{G}$ on $\hat{\Omega}$ such that the cocycle $\hat{c}:H\times X_{\hat{G}}\to\hat{G}$ is $L^{1}$-integrable. Let $S=\{g_{1},\dots,g_{k}\}$ be a (finite) set of representatives of the right cosets of $G$ in $\hat{G}$. Then $X_{G}=g_{1}X_{\hat{G}}\cup\dots\cup g_{k}X_{\hat{G}}$ is a fundamental domain for the $G$-action on $\hat{\Omega}$. We claim that the associated cocycle $c:H\times X_{G}\to G$ is $L^{1}$-integrable. Indeed, let $h\in H$. For $x\in X_{G}$, there exists a unique $\hat{x}\in X_{\hat{G}}$ and $j\in\{1,\dots,k\}$ such that $x=g_{j}\hat{x}$. Also by definition of $X_{\hat{G}}$, there exists a unique element $\hat{g}\in\hat{G}$ such that $\hat{g}hx\in X_{\hat{G}}$. Using the fact that the actions of $\hat{G}$ and $H$ on $\hat{\Omega}$ commute, we see that $(\hat{g}g_{j})h\hat{x}\in X_{\hat{G}}$, showing that $\hat{c}(h,\hat{x})=\hat{g}g_{j}$. On the other hand, the set $S^{-1}=\{g_{1}^{-1},\dots,g_{k}^{-1}\}$ is a set of representatives of the left cosets of $G$ in $\hat{G}$, so there exist $i\in\{1,\dots,k\}$ and $g\in G$ such that $\hat{g}=g_{i}^{-1}g$. We then have $ghx\in g_{i}X_{\hat{G}}$, showing that $c(h,x)=g$. It follows that $c(h,x)\in S\hat{c}(h,\hat{x})S^{-1}$. As this is true for every $h\in H$ and $x\in X_{G}$ and $S$ is a finite set, the lemma follows. ∎
@@
Our proof of Lemma 7.3 is inspired by an argument coming from ongoing work of Escalier and the first-named author [EH].
@@
As above, let $\hat{X}_{G}$ be a fundamental domain for the action of $G$ on $\hat{\Omega}$, such that the cocycle $\hat{c}:H\times\hat{X}_{G}\to G$ is $L^{1}$-integrable. By Lemma 6.7, the space $\hat{\Omega}_{v}=\theta^{-1}(\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v))$ is a measure equivalence coupling between $G_{v}$ and $H_{v}$. Let $Y_{G_{v}},Y_{H_{v}}$ be respective Borel fundamental domains for these actions on $\hat{\Omega}_{v}$. Lemma 6.7 also shows that $Y_{G_{v}},Y_{H_{v}}$ extend to Borel fundamental domains $Y_{G},Y_{H}$ for the actions of $G$ and $H$ on $\hat{\Omega}$. Let $c_{v}:H_{v}\times Y_{G_{v}}\to G_{v}$ be the measure equivalence cocycle. Then $c_{v}$ extends to a measure equivalence cocycle $c^{\prime}:H\times Y_{G}\to G$. The cocycle $c^{\prime}$ is cohomologous to $\hat{c}$, i.e. there exists $\theta:Y_{G}\to G$ such that $c^{\prime}(h,x)=\theta(h\cdot x)^{-1}\hat{c}(h,\tilde{x})\theta(x)$, where $\tilde{x}$ is the unique element of $\hat{X}_{G}$ in the same $G$-orbit as $x$. As usual $G$ is equipped with its standard generating set. Let $\psi:G\to G$ be defined by letting $\psi(g)$ be the unique element with smallest word length such that $\psi(g)G_{v}=gG_{v}$ (uniqueness comes from the normal form in the right-angled Artin group, in the sense of graph products [Gre90]). Let $\theta^{\prime}=\psi\circ\theta$, and let $c^{\prime\prime}:H\times Y_{G}\to G$ be defined by $c^{\prime\prime}(h,x)=\theta^{\prime}(h\cdot x)^{-1}\hat{c}(h,\tilde{x})\theta^{\prime}(x)$. By definition of $\psi$, for every $x\in X$, there exists $g_{x}\in G_{v}$ such that $\theta^{\prime}(x)=\theta(x)g_{x}$. Using this and the fact that $c^{\prime}(H_{v}\times Y_{G_{v}})\subseteq G_{v}$, we deduce that $c^{\prime\prime}(H_{v}\times Y_{G_{v}})\subseteq G_{v}$. In addition, a normal form argument shows that $|c^{\prime\prime}(h,x)|_{G_{v}}\leq|\hat{c}(h,\tilde{x})|_{G}$ for every $h\in H_{v}$ and $x\in Y_{G_{v}}$ (indeed, one writes $\theta^{\prime}(hx)c^{\prime\prime}(h,x)\theta^{\prime}(x)^{-1}=\hat{c}(h,\tilde{x})$, and observes that the subword $c^{\prime\prime}(h,x)$ on the left cannot be shortened by cancellations with $\theta^{\prime}(x)$ and $\theta^{\prime}(hx)$ in view of the choice of $\theta^{\prime}$). Therefore the cocycle $c^{\prime\prime}:H_{v}\times Y_{G_{v}}\to G_{v}$ is $L^{1}$-integrable. It is $G_{v}$-cohomologous to $c^{\prime}$, and therefore it is a measure equivalence cocycle from $H_{v}$ to $G_{v}$, for the coupling $\Omega_{v}$. This completes our proof. ∎
@@ 7.2 Integrable embeddings
A cocycle $c:H\times X\to G$ is an $L^{1}$-integrable embedding if for every $\varepsilon>0$, there exist an $L^{1}$-integrable cocycle $c^{\prime}:H\times X\to G$ which is cohomologous to $c$, a Borel subset $X_{0}\subseteq X$ with $\mu(X_{0})\geq 1-\varepsilon$, and a constant $C=C(\varepsilon)$, such that for every $h\in H$ and almost every $x\in X_{0}$, one has |$|\{g\in G\mid gx\in X_{0}\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }c^{\prime}(g,x)=h\}|<C.$
@@
Let $\Omega$ be a measure equivalence coupling between two countable groups $G$ and $H$, let $X_{G}\subseteq\Omega$ be a fundamental domain for the $H$-action, and let $c:H\times X_{G}\to G$ be the associated measure equivalence cocycle. Let $\varepsilon>0$. Then there exist $C>0$ and a measurable subset $X_{\varepsilon}$ of $X_{G}$ with $\mu(X_{\varepsilon})\geq\mu(X_{G})-\varepsilon$ such that for every $x\in X_{\varepsilon}$ and every $g\in G$, one has
|$|\{h\in H\mid h\cdot x\in X_{\varepsilon}\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }c(h,x)=g\}|<C.$
@@
Let $X_{H}$ be a fundamental domain for the $H$-action on $\Omega$. Then $\Omega$ is covered by countably many pairwise disjoint $H$-translates of $X_{H}$. For every $h\in H$, let $X_{G,h}=X_{G}\cap hX_{H}$. Then the subsets $X_{G,h}$ form a countable partition of $X_{G}$. Choose a finite subset $F\subseteq H$ such that $\mu(\cup_{h\in F}X_{G,h})>\mu(X_{G})-\varepsilon$. Let $C=|F|$, and let $X_{\varepsilon}=\cup_{g\in F}X_{G,h}$. Let now $g\in G$ and $x\in X_{\varepsilon}$. We claim that if two distinct elements $h_{1},h_{2}\in H$ satisfy $c(h_{1},x)=c(h_{2},x)=g$, then $h_{1}x$ and $h_{2}x$ belong to different subsets of the partition $g^{-1}X_{G}=\sqcup g^{-1}X_{G,h}$. Indeed, notice first that as $c(h_{1},x)=c(h_{2},x)=g$, we have $h_{1}x,h_{2}x\in g^{-1}X_{G}$. Arguing by contradiction that the claim fails, there exists $h\in H$ such that $h_{1}x,h_{2}x\in g^{-1}(X_{G}\cap hX_{H})$. In particular $gh_{1}x,gh_{2}x\in hX_{H}$. As the actions of $G$ and $H$ on $\Omega$ commute, it follows that $h_{1}$ and $h_{2}$ both send $gx$ to a common fundamental domain (namely $hX_{H}$) of the $H$-action on $\Omega$. This is a contradiction, which proves the claim.
@@ 7.3 Stabilizers of rank $1$ vertices are virtually cyclic
In this section, we prove Proposition 7.1. We already know that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H_{v}$ to $G_{v}$ (Lemma 7.3). If we additionally knew that $H_{v}$ were finitely generated, then we could use a theorem of Bowen [Aus16b, Theorem B.10] to deduce that the growth of $H_{v}$ is at most equal to the growth of $G_{v}$, whence at most linear. Being infinite, $H_{v}$ is therefore virtually isomorphic to $\mathbb{Z}$. However, we do not know a priori that $H_{v}$ is finitely generated. In this section, we extend Bowen’s theorem in order to deal with this issue, see Theorem 7.10. Given a finitely generated group $G$, with a finite generating set $S$, we let $B_{G,S}(e,n)=\{g\in G\mid|g|_{S}\leq n\}$. The growth function $\mathrm{gr}_{G,S}:\mathbb{N}\to\mathbb{N}$ is defined as
@@
Let $L$ be a countable group, let $G$ be a finitely generated group, and assume that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $L$ to $G$. Then for every finitely generated subgroup $L^{\prime}$ of $L$, one has $\mathrm{gr}_{L^{\prime}}\preceq\mathrm{gr}_{G}$.
@@
[…]
@@
[…]
@@
Recall that a group $L$ is locally virtually cyclic if every finitely generated subgroup of $L$ is virtually cyclic (possibly finite).
Lemma 7.9.
Let $L$ be a locally virtually cyclic countable group with bounded torsion. Then either $L$ is virtually cyclic, or else there exists an infinite cyclic subgroup $L_{1}\subseteq L$ such that for every $k\in\mathbb{N}$, there exists an infinite cyclic subgroup $L_{k}\subseteq L$ that contains $L_{1}$ as a subgroup of index at least $k$.
@@
Suppose first that $L$ is locally finite, i.e. every finitely generated subgroup is finite. Then $L$ must be finite because of the bounded torsion assumption.
Now assume $L$ is not locally finite. The conclusion is obvious if $L$ is finitely generated, so we assume otherwise. The group $L$ has an infinite finitely generated subgroup, which is virtually cyclic by our assumption. Thus there exists an infinite cyclic subgroup $L^{\prime}_{1}\subseteq L$. By adding extra elements successively, we can find an infinite increasing sequence of finitely generated, whence virtually cyclic, subgroups $L^{\prime}_{1}\subsetneq L^{\prime}_{2}\subsetneq\dots$ In particular, as $k$ goes to +$\infty$+ ∞, the index $[L^{\prime}_{k}:L^{\prime}_{1}]$ goes to +$\infty$+ ∞. Let $K$ be a bound on the cardinality of a finite subgroup of $L$. Observe that every infinite virtually cyclic group with no finite subgroup of cardinality larger than $K$, contains an infinite cyclic subgroup of index at most $2K$: this follows, for instance, from the fact that every infinite virtually cyclic group is either $-\mathbb{Z}$ or finite-by-dihedral. For every $k\in\mathbb{N}$, let $L^{\prime\prime}_{k}\subseteq L^{\prime}_{k}$ be an infinite cyclic subgroup of order at most $2K$. Letting $L_{1}(k)=L^{\prime}_{1}\cap L^{\prime\prime}_{k}$, there is a bound on the index $[L^{\prime}_{1}:L_{1}(k)]$. Therefore, up to a subsequence, we can assume that all groups $L_{1}(k)$ are equal to the same group $L_{1}$. As $L_{1}\subseteq L^{\prime\prime}_{k}$ and the index $[L^{\prime\prime}_{k}:L_{1}]$ goes to +$\infty$+ ∞, the lemma follows (by letting $L_{k}=L^{\prime\prime}_{\sigma(k)}$ for some increasing map $\sigma:\mathbb{N}\to\mathbb{N}$ chosen so that $[L^{\prime\prime}_{\sigma(k)}:L_{1}]\geq k$). ∎
@@
Let $L$ be a countable group with bounded torsion, and assume that there exists an $L^{1}$-integrable embedding from $L$ to $\mathbb{Z}$. Then $L$ is virtually cyclic (possibly finite).
@@
We write $J=\mathbb{Z}$. We choose once and for all a generator for $J$, and consider the associated word length on $J$ for the rest of the argument. For $C>0$, we say that a map between two sets $f:E_{1}\to E_{2}$ is at most $C$-to-1 if the cardinality of each point inverse of $f$ is at most $C$. Without loss of generality, we can assume that $L$ is infinite. By Lemma 7.7, every finitely generated subgroup of $L$ has growth at most linear, and therefore is virtually cyclic (possibly finite), see e.g. [Jus71]. Assuming towards a contradiction that $L$ is not virtually cyclic, let $L_{1}\subseteq L$ be an infinite cyclic subgroup of $L$ given by Lemma 7.9. We will prove that there exists $K>0$ such that for every other infinite cyclic subgroup $L_{2}$ containing $L_{1}$, one has $[L_{2}:L_{1}]\leq K$. This will contradict Lemma 7.9 and conclude our proof.
By assumption, there exists a measure-preserving action of $L$ on a standard probability space $(X,\mu)$, a cocycle $c:L\times X\to J$ which is $L^{1}$-integrable, a Borel subset $X_{0}\subseteq X$ with $\mu(X_{0})\geq 0.9$, and a constant $C>0$, such that for every $x\in X_{0}$, the restriction $c(\cdot,x)_{|R_{X_{0}}(x)}$ is at most $C$-to-$1$. Let $L_{2}$ be an infinite cyclic subgroup of $L$ that contains $L_{1}$, and let $k=[L_{2}:L_{1}]$. For every $i\in\{1,2\}$, we choose a generator $g_{i}$ of $L_{i}$, and we endow $L_{i}$ with the word metric with respect to this choice of generator. For $g\in L_{2}$, let
[…]
Let $M_{1}=\kappa(g_{1})$ and $D=\max\{\kappa(g_{2}^{i})\}_{0\leq i\leq k-1}$. Let $g\in L_{2}$, and write it as $g=\ell g_{1}+ig_{2}$ with $\ell\in\mathbb{Z}$ and $10\leq i\leq k-$ 1. Using the cocycle relation, we see that $\kappa(g)\leq|\ell|M_{1}+D$. On the other hand $|g|_{L_{2}}=k|\ell|\pm i$. Therefore, there exists $M^{\prime}\in\mathbb{N}$ such that $\kappa(g)\leq\frac{2M_{1}}{k}|g|_{L_{2}}$ whenever $|g|_{L_{2}}\geq M^{\prime}$. Recall that $\mu(X_{0})\geq 0.9$. We can therefore apply Lemma 7.8 with $M=\frac{2M_{1}}{k}$, and with $L_{2}$ in place of $L$, and deduce that for $n$ large enough, there exists a Borel subset $Y_{n}\subseteq X_{0}$ with $\mu(Y_{n})\geq 0.3$ such that for every $x\in Y_{n}$, one has

### v2

@@ 7 Exploiting integrability
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $\mathbb{B}$ be its right-angled building, and let $H$ be a countable group with bounded finite subgroups. Assume that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$.
Then $H$ acts on $\mathbb{B}$ with virtually infinite cyclic stabilizers of rank $1$ vertices.
The plan of the proof is to show first that for each vertex $v\in V\mathbb{B}$ there is an $(L^{1},L^{0})$-measure equivalence coupling from $H_{v}$ (the $H$-stabilizer of $v$) to $G_{v}$ (Lemma 7.3), which then gives an $L^{1}$-integrable embedding from $H_{v}$ to $G_{v}$ (Corollary 7.6). When $v$ is rank 1, we have $G_{v}\cong\mathbb{Z}$. Then we show in Theorem 7.11 that $H_{v}$ is virtually $\mathbb{Z}$: if we knew that $H_{v}$ were finitely generated, then this would follow from work of Bowen (in an appendix of an article by Austin [Aus16b, Theorem B.10]) saying that growth is invariant under $L^{1}$-measure equivalence; the main point of our work is to extend this to the case where $H_{v}$ is possible infinitely generated.
@@ 7.1 Integrable coupling between vertex stabilizers
Let $G,H$ be countable groups, with $G$ finitely generated, and let $\hat{G}$ be a finite-index extension of $G$. Let $\hat{\Omega}$ be an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $\hat{G}$.
Then $\hat{\Omega}$ is an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$.
@@
By definition, there exists a fundamental domain $X_{\hat{G}}$ for the action of $\hat{G}$ on $\hat{\Omega}$ such that the cocycle $\hat{c}:H\times X_{\hat{G}}\to\hat{G}$ is $L^{1}$-integrable.
Let $S=\{g_{1},\dots,g_{k}\}$ be a (finite) set of representatives of the right cosets of $G$ in $\hat{G}$. Then $X_{G}=g_{1}X_{\hat{G}}\cup\dots\cup g_{k}X_{\hat{G}}$ is a fundamental domain for the $G$-action on $\hat{\Omega}$.
We claim that the associated cocycle $c:H\times X_{G}\to G$ is $L^{1}$-integrable. Indeed, let $h\in H$. For $x\in X_{G}$, there exists a unique $\hat{x}\in X_{\hat{G}}$ and $j\in\{1,\dots,k\}$ such that $x=g_{j}\hat{x}$. Also by definition of $X_{\hat{G}}$, there exists a unique element $\hat{g}\in\hat{G}$ such that $\hat{g}hx\in X_{\hat{G}}$. Using the fact that the actions of $\hat{G}$ and $H$ on $\hat{\Omega}$ commute, we see that $(\hat{g}g_{j})h\hat{x}\in X_{\hat{G}}$, showing that $\hat{c}(h,\hat{x})=\hat{g}g_{j}$. On the other hand, the set $S^{-1}=\{g_{1}^{-1},\dots,g_{k}^{-1}\}$ is a set of representatives of the left cosets of $G$ in $\hat{G}$, so there exist $i\in\{1,\dots,k\}$ and $g\in G$ such that $\hat{g}=g_{i}^{-1}g$. We then have $ghx\in g_{i}X_{\hat{G}}$, showing that $c(h,x)=g$. It follows that $c(h,x)\in S\hat{c}(h,\hat{x})S^{-1}$. As this is true for every $h\in H$ and $x\in X_{G}$ and $S$ is a finite set, the lemma follows. ∎
@@
Our proof of Lemma 7.3 is inspired by an argument coming from work of Escalier and the first-named author [EH24, Theorem 12.1].
@@
As above, let $\hat{X}_{G}$ be a fundamental domain for the action of $G$ on $\hat{\Omega}$, such that the cocycle $\hat{c}:H\times\hat{X}_{G}\to G$ is $L^{1}$-integrable.
By Lemma 6.7, the space $\hat{\Omega}_{v}=\theta^{-1}(\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v))$ is a measure equivalence coupling between $G_{v}$ and $H_{v}$. Let $Y_{G_{v}},Y_{H_{v}}$ be respective Borel fundamental domains for these actions on $\hat{\Omega}_{v}$. Lemma 6.7 also shows that $Y_{G_{v}},Y_{H_{v}}$ extend to Borel fundamental domains $Y_{G},Y_{H}$ for the actions of $G$ and $H$ on $\hat{\Omega}$.
Let $c_{v}:H_{v}\times Y_{G_{v}}\to G_{v}$ be the measure equivalence cocycle. Then $c_{v}$ extends to a measure equivalence cocycle $c^{\prime}:H\times Y_{G}\to G$. The cocycle $c^{\prime}$ is cohomologous to $\hat{c}$, i.e. there exists $\theta:Y_{G}\to G$ such that $c^{\prime}(h,x)=\theta(h\cdot x)^{-1}\hat{c}(h,\tilde{x})\theta(x)$, where $\tilde{x}$ is the unique element of $\hat{X}_{G}$ in the same $G$-orbit as $x$.
As usual $G$ is equipped with its standard generating set. Let $\psi:G\to G$ be defined by letting $\psi(g)$ be the unique element with smallest word length such that $\psi(g)G_{v}=gG_{v}$ (uniqueness comes from the normal form in the right-angled Artin group, in the sense of graph products [Gre90]).
[…]
Therefore the cocycle $c^{\prime\prime}:H_{v}\times Y_{G_{v}}\to G_{v}$ is $L^{1}$-integrable. It is $G_{v}$-cohomologous to $c^{\prime}$, and therefore it is a measure equivalence cocycle from $H_{v}$ to $G_{v}$, for the coupling $\Omega_{v}$. This completes our proof. ∎
@@ 7.2 Integrable embeddings
A cocycle $c:H\times X\to G$ is an $L^{1}$-integrable embedding if for every $\varepsilon>0$, there exist an $L^{1}$-integrable cocycle $c^{\prime}:H\times X\to G$ which is cohomologous to $c$, a Borel subset $X_{0}\subseteq X$ with $\mu(X_{0})\geq 1-\varepsilon$, and a constant $C=C(\varepsilon)$, such that for every $h\in H$ and almost every $x\in X_{0}$, one has $|\{g\in G\mid gx\in X_{0}\text{\penalty 10000\ and\penalty 10000\ }c^{\prime}(g,x)=h\}|<C.$
@@
Let $\Omega$ be a measure equivalence coupling between two countable groups $G$ and $H$, let $X_{G}\subseteq\Omega$ be a fundamental domain for the $H$-action, and let $c:H\times X_{G}\to G$ be the associated measure equivalence cocycle. Let $\varepsilon>0$.
Then there exist $C>0$ and a measurable subset $X_{\varepsilon}$ of $X_{G}$ with $\mu(X_{\varepsilon})\geq\mu(X_{G})-\varepsilon$ such that for every $x\in X_{\varepsilon}$ and every $g\in G$, one has
$|\{h\in H\mid h\cdot x\in X_{\varepsilon}\text{\penalty 10000\ and\penalty 10000\ }c(h,x)=g\}|<C.$
@@
Let $X_{H}$ be a fundamental domain for the $H$-action on $\Omega$. Then $\Omega$ is covered by countably many pairwise disjoint $H$-translates of $X_{H}$. For every $h\in H$, let $X_{G,h}=X_{G}\cap hX_{H}$. Then the subsets $X_{G,h}$ form a countable partition of $X_{G}$. Choose a finite subset $F\subseteq H$ such that $\mu(\cup_{h\in F}X_{G,h})>\mu(X_{G})-\varepsilon$. Let $C=|F|$, and let $X_{\varepsilon}=\cup_{g\in F}X_{G,h}$.
Let now $g\in G$ and $x\in X_{\varepsilon}$. We claim that if two distinct elements $h_{1},h_{2}\in H$ satisfy $c(h_{1},x)=c(h_{2},x)=g$, then $h_{1}x$ and $h_{2}x$ belong to different subsets of the partition $g^{-1}X_{G}=\sqcup g^{-1}X_{G,h}$. Indeed, notice first that as $c(h_{1},x)=c(h_{2},x)=g$, we have $h_{1}x,h_{2}x\in g^{-1}X_{G}$. Arguing by contradiction that the claim fails, there exists $h\in H$ such that $h_{1}x,h_{2}x\in g^{-1}(X_{G}\cap hX_{H})$. In particular $gh_{1}x,gh_{2}x\in hX_{H}$. As the actions of $G$ and $H$ on $\Omega$ commute, it follows that $h_{1}$ and $h_{2}$ both send $gx$ to a common fundamental domain (namely $hX_{H}$) of the $H$-action on $\Omega$. This is a contradiction, which proves the claim.
@@ 7.3 Stabilizers of rank $1$ vertices are virtually cyclic
In this section, we prove Proposition 7.1. We already know that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H_{v}$ to $G_{v}$ (Lemma 7.3). If we additionally knew that $H_{v}$ were finitely generated, then we could use a theorem of Bowen [Aus16b, Theorem B.10] to deduce that the growth of $H_{v}$ is at most equal to the growth of $G_{v}$, whence at most linear. Being infinite, $H_{v}$ is therefore virtually isomorphic to $\mathbb{Z}$. However, we do not know a priori that $H_{v}$ is finitely generated. In this section, we extend Bowen’s theorem in order to deal with this issue, see Theorem 7.11.
Given a finitely generated group $G$, with a finite generating set $S$, we let $B_{G,S}(e,n)=\{g\in G\mid|g|_{S}\leq n\}$. The growth function $\mathrm{gr}_{G,S}:\mathbb{N}\to\mathbb{N}$ is defined as
@@
Let $L$ be a countable group, let $G$ be a finitely generated group, and assume that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $L$ to $G$.
Then for every finitely generated subgroup $L^{\prime}$ of $L$, one has $\mathrm{gr}_{L^{\prime}}\preceq\mathrm{gr}_{G}$.
@@
Let $L$ and $J$ be finitely generated groups, with $L$ infinite, equipped with word metrics $|\cdot|_{L}$ and $|\cdot|_{J}$ (associated to finite generating sets). Let $L\curvearrowright(X,\mu)$ be a measure-preserving action on a standard probability space, and $c:L\times X\to J$ an $L^{1}$-integrable cocycle. Let $M^{\prime}>0$ and $M>0$ be such that for every $g\in L$ satisfying $|g|_{L}\geq M^{\prime}$, one has $\int_{X}|c(g,x)|_{J}\;d\mu(x)\leq M|g|_{L}$. Let $X_{0}\subseteq X$ be a measurable subset with $\mu(X_{0})\geq 0.9$.
Then there exists $n_{0}>0$ such that for every $n\geq n_{0}$, there exists a measurable subset $Y_{n}\subset X_{0}$ with $\mu(Y_{n})\geq 0.3$, such that for every $x\in Y_{n}$, one has
@@
The existence of $M^{\prime}$ and $M$ follows from the cocycle identity and subadditivity, see e.g. [BFS13, Section A.2].
@@
We follow Bowen’s proof of [Aus16b, Theorem B.10] very closely. For $g\in L$, we write $\kappa(g)=\displaystyle\int_{X}|c(g,x)|_{J}\;d\mu(x)$. As $L$ is infinite and there are only finitely many elements in $L$ such that $\kappa(g)>M|g|_{L}$, there exists $n_{0}$ such that for every $n\geq n_{0}$, one has
@@
Recall that a group $L$ is locally virtually cyclic if every finitely generated subgroup of $L$ is virtually cyclic (possibly finite). We are grateful to Yves Cornulier for providing us with the following lemma and its proof, improving a lemma we were using in an earlier version of this paper.
Lemma 7.9 (Cornulier).
Let $L$ be a locally virtually cyclic countable group with bounded finite subgroups. Then $L$ has a characteristic subgroup of finite-index which is torsion-free and locally cyclic.
@@
We will assume that $L$ is infinite, as otherwise the conclusion is obvious. We will take advantage of a result of Wall [Wal67, Lemma 4.1], ensuring that every group which is virtually isomorphic to $\mathbb{Z}$ surjects onto either $\mathbb{Z}$ or onto the infinite dihedral group $D_{\infty}$ with finite kernel. This justifies the case disjunction made in the next two paragraphs.
We first assume that every infinite finitely generated subgroup of $L$ surjects onto $\mathbb{Z}$ with finite kernel. In this case the set $F$ of finite-order elements of $L$ is a subgroup, which by assumption is finite (and it is characteristic). The quotient $L/F$ is torsion-free, so every finitely generated subgroup of $L/F$ is isomorphic to $\mathbb{Z}$. Let $M\subseteq L$ be the centralizer of $F$. Then $M$ has finite index in $L$, it is characteristic, and it is central-by-(locally cyclic) hence abelian. Let $n$ be the order of $F\cap M$. Then the subgroup $M^{n}$ consisting of $n^{\text{th}}$ powers of elements of $M$ intersects $F$ trivially, and has finite index, so is the desired locally cyclic finite-index characteristic subgroup of $M$.
We now assume that $L$ contains a subgroup $H$ that maps to $D_{\infty}$ with finite kernel. Let $\theta_{H}:H\to\mathbb{Z}/2\mathbb{Z}$ be the homomorphism given by the action of $H$ on its two ends. Every finitely generated subgroup $G$ of $L$ containing $H$ must then also surjects onto $D_{\infty}$ with finite kernel, and the homomorphism $\theta_{G}$ extends $\theta_{H}$. Altogether these give a homomorphism $L\to\mathbb{Z}/2\mathbb{Z}$, whose kernel is a characteristic finite-index subgroup to which the above paragraph applies. ∎
We record the following corollary, which is the form in which Lemma 7.9 will be applied in the sequel.
Corollary 7.10.
Let $L$ be a locally virtually cyclic countable group with bounded finite subgroups. Then either $L$ is virtually cyclic, or else there exists an infinite cyclic subgroup $L_{1}\subseteq L$ such that for every $k\in\mathbb{N}$, there exists an infinite cyclic subgroup $L_{k}\subseteq L$ that contains $L_{1}$ as a subgroup of index at least $k$.
Proof.
Let $L^{0}$ be a torsion-free locally cyclic finite-index subgroup of $L$: this exists by Lemma 7.9. If $L^{0}$ is finitely generated, then $L$ is virtually cyclic. Otherwise, we can find an infinite strictly increasing sequence $L_{1}\subseteq L_{2}\subseteq\dots$ of finitely generated subgroups of $L^{0}$, whence all isomorphic to $\mathbb{Z}$. In particular $L_{k}$ contains $L_{1}$ as a subgroup of index at least $k$ (in fact at least $2^{k-1}$), and we are done. ∎
@@
Let $L$ be a countable group with bounded finite subgroups, and assume that there exists an $L^{1}$-integrable embedding from $L$ to $\mathbb{Z}$.
Then $L$ is virtually cyclic (possibly finite).
@@
We write $J=\mathbb{Z}$. We choose once and for all a generator for $J$, and consider the associated word length on $J$ for the rest of the argument. For $C>0$, we say that a map between two sets $f:E_{1}\to E_{2}$ is at most $C$-to-1 if the cardinality of each point inverse of $f$ is at most $C$.
Without loss of generality, we can assume that $L$ is infinite. By Lemma 7.7, every finitely generated subgroup of $L$ has growth at most linear, and therefore is virtually cyclic (possibly finite), see e.g. [Jus71]. Assuming towards a contradiction that $L$ is not virtually cyclic, let $L_{1}\subseteq L$ be an infinite cyclic subgroup of $L$ given by Corollary 7.10. We will prove that there exists $K>0$ such that for every other infinite cyclic subgroup $L_{2}$ containing $L_{1}$, one has $L_{2}:L_{1}]\leq K$. This will contradict Corollary 7.10 and conclude our proof.
By assumption, there exists a measure-preserving action of $L$ on a standard probability space $(X,\mu)$, a cocycle $c:L\times X\to J$ which is $L^{1}$-integrable, a Borel subset $X_{0}\subseteq X$ with $\mu(X_{0})\geq 0.9$, and a constant $C>0$, such that for every $x\in X_{0}$, the restriction $c(\cdot,x)_{|R_{X_{0}}(x)}$ is at most $C$-to-11.
Let $L_{2}$ be an infinite cyclic subgroup of $L$ that contains $L_{1}$, and let $k=[L_{2}:L_{1}]$. For every $i\in\{1,2\}$, we choose a generator $g_{i}$ of $L_{i}$, and we endow $L_{i}$ with the word metric with respect to this choice of generator. For $g\in L_{2}$, let
$\kappa(g)=\int_{X}|c(g,x)|_{J}\;d\mu(x).$
Let $M_{1}=\kappa(g_{1})$ and $D=\max\{\kappa(g_{2}^{i})\}_{0\leq i\leq k-1}$. Let $g\in L_{2}$, and write it as $g=\ell g_{1}+ig_{2}$ with $\ell\in\mathbb{Z}$ and $0\leq i\leq k-1$. Using the cocycle relation, we see that $\kappa(g)\leq|\ell|M_{1}+D$. On the other hand $|g|_{L_{2}}=k|\ell|\pm i$. Therefore, there exists $M^{\prime}\in\mathbb{N}$ such that $\kappa(g)\leq\frac{2M_{1}}{k}|g|_{L_{2}}$ whenever $|g|_{L_{2}}\geq M^{\prime}$.
Recall that $\mu(X_{0})\geq 0.9$. We can therefore apply Lemma 7.8 with $M=\frac{2M_{1}}{k}$, and with $L_{2}$ in place of $L$, and deduce that for $n$ large enough, there exists a Borel subset $Y_{n}\subseteq X_{0}$ with $\mu(Y_{n})\geq 0.3$ such that for every $x\in Y_{n}$, one has